OpenStudy (anonymous):

snowys snow comes has a special bubble gum snow cone on sale. the cone is regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. the radius of the snow cone is 2 inches and the height of the cone is 3 inches. is the diameter of the bubble gum is 0.5 inches , what is the closest approximation of the volume of the cone that can be filled with flavored ice. use 3.14 as pi a- 22.97 cubic in b-12.50 cubic in c-13.24 cubic in d-12.60 cubic in medal and fan will be given

OpenStudy (anonymous):

- isnt negative btw

OpenStudy (anonymous):

@jdoe0001 hi can you help me please

OpenStudy (anonymous):

@campbell_st hi can you help me again please :)

OpenStudy (anonymous):

@ganeshie8 can you help me please

OpenStudy (anonymous):

@mathmale hey can you help me please

OpenStudy (anonymous):

@IMStuck @jdoe0001 help please

OpenStudy (campbell_st):

You need volume of the cone - volume of the sphere \[V = \frac{1}{3} \times 3.14 \times 2 \times 3 - \frac{4}{3} \times 3.14 \times 0.5^3\]